OpenStudy (anonymous):

The equation 2x^2 + x -4 =0 has roots a and beta and the equation x^2 - 2x + p =0 has roots k(a)/beta and k(beta)/a. Find the value of k and of p.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

first find what a and beta are I think

OpenStudy (phi):

2x^2 + x -4 =0 has roots a and beta means you can factor the quadratic into 2(x-a)(x-b)=0 expanding, you get 2(x^2 - a x - b x +ab) =0 or 2x^2 - 2(a+b) x + 2ab = 0 match corresponding terms to find: -2(a+b) = 1 or (a+b)= -1/2 2ab= -4 or ab = -2 now do the same thing for the 2nd polynomial x^2 - 2x + p =0 has roots k(a)/beta and k(beta)/a expand (x - k a/b)(x - k b/a) and match corresponding terms of x^2 - 2x + p =0 you will need to use the "trick" that (a+b)^2 = a^2 + b^2 +2ab from which you get a^2 + b^2 = (a+b)^2 - 2ab